• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name adaq

    Log entry time 00:03:03 on April 25, 2001

    Entry number 61302

    keyword=Beam Time Accounting
    Hall A

    4/24/2001 1600-2400
    Experiment #: commissioning
    Shift Leader: P. King
    Shift Workers: A. Camsonne, P. Roos

    16001700180019002000210022002300Total Time
    Acceptable Beam
    in Use
    0 5 15 50 40 45 40 35 3.83333333333333
    Beam Available
    but Not in Use
    1 55 0 10 20 15 20 20 3.33333333333333
    Beam Not Available
    or Unacceptable
    0 0 45 0 0 0 0 5 0.833333333333333
    Accelerator Con-
    figuration Chage
    0 0 0 0 0 0 0 0 0
    Experiment
    Ready
    0 5 50 50 40 45 40 40 4.5
    Planned Config-
    uration Change
    1 55 10 10 20 15 20 20 3.5
    Unplanned
    Experiment Down
    0 0 0 0 0 0 0 00
    Hall not
    ready reasons
    Target 0 0 0 0 0 0 0 0 0
    HRSe 0 0 0 0 0 0 0 0 0
    HRSh 0 0 0 0 0 0 0 0 0
    Data Ac 0 0 0 0 0 0 0 0 0
    Other 0 0 0 0 0 0 0 0 0

    Comments: 16:55 Target ready for beam. 17:55 Right dipole reaches momentum setpoint. 18:14-18:24 Change collimator settings 19:25-19:35 Change spectrometer momentum. 20:00-20:08 Change collimation setting. 20:28-20:38 Change spectrometer momentum. 20:58-21:02 Move to the optics target. 21:10-21:17 Change collimation. 21:29-21:33 Change to the LH2 target. 22:00-22:08 Change spectrometer momenta. 22:31-22:37 Change spectrometer collimation. 22:54-23:04 Change to the optics target and change spectrometer momenta. 23:09-23:15 Change spectrometer collimation. 23:20-23:25 Change to the LH2 target. 23:50-23:55 Change spectrometer collimation. 23:55 The injector is having problems.

    Data Entry by: P. King